OpenStudy (anonymous):

Write a quadratic function that fits the given set of points. (2, 3), (6, 3), and (8, -3)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I believe you did exactly the right thing before. Is this the same problem as before, or a new one? I thought we were at the point where we needed to solve a system of linear equations.

OpenStudy (anonymous):

The same one!

OpenStudy (mathmale):

Here's where you were: 3 = 4a + 2b + c 3 = 36a + 6b + c -3 = 64a + 8b + c We were talking about how to solve this system for the constants a, b and c. I asked you to identify the method you'd prefer to use to do so. Just say what you want to do, and we'll go ahead with it.

OpenStudy (mathmale):

Here are some of the methods you have to choose from: elimination by addition and subtraction elim. by substitution matrix algebra determinants

OpenStudy (anonymous):

I am sorry that it has take so long to reply. My girlfriend called to cheer me up, I have been stressed today, it has been a busy day with church bands and work. But I am ready to get this work done! The only one of those options I am really familiar with is Elimination by subtraction.

OpenStudy (mathmale):

We'll do the problem that way then. I'm going to modify your set of equations just a little: from: 3 = 4a + 2b + c 3 = 36a + 6b + c -3 = 64a + 8b + c 4a + 2b + c = 3 36a +6b +c =3 64a+8b+c=-3

OpenStudy (mathmale):

All i did was to move the constants to the right side from the left side. Look at the first 2 equations, Rick: How would you go about eliminating c?

OpenStudy (mathmale):

36a +6b +c =3 -(4a + 2b + c = 3) ----------------- ?

OpenStudy (anonymous):

Would I subtract one from the other? 4a + 2b + c = 3 - 36a +6b +c =3 ---------------

OpenStudy (anonymous):

Oh yes! Ha! I was right!

OpenStudy (mathmale):

Yes. Please go ahead with it now.

OpenStudy (mathmale):

Suggest doing it on paper if you aren't already. Helps when you can see what you've already done. You might want to save this paper and keep such papers in order so that you have something to review later.

OpenStudy (anonymous):

Yes, I am doing it on paper, I will have it finished in just a bit.

OpenStudy (anonymous):

It would be: -32a - 4b = 0

OpenStudy (mathmale):

That's great. I have the same result, except I chose to use all + signs. Nice work! Now do exactly the same thing with the 1st and 3rd equations; elim c again. Hold on to your first result; we'll use it later.

OpenStudy (anonymous):

So. 4a + 2b + c = 3 - 64a + 8b + c = -3 ------------------ ?

OpenStudy (mathmale):

Yes. try it.

OpenStudy (anonymous):

-60a - 6b = 6 ??

OpenStudy (mathmale):

Again, perfect! You do know your stuff. Now y ou have 2 equations in 2 unknowns, with c gone.

OpenStudy (anonymous):

Yes!

OpenStudy (mathmale):

-60a - 6b = 6 -32a - 4b = 0

OpenStudy (mathmale):

How would you go about eliminating either a or b (your choice)?

OpenStudy (anonymous):

Divide from both sides.

OpenStudy (anonymous):

or would I subtract from bot sides?

OpenStudy (mathmale):

Unsure of what you mean. What would the purpose of doing that be? I do see that both equations can be reduced; is that what you meant?

OpenStudy (anonymous):

In order to isolate "a", would I subtract -32 or -60 from both sides of the equation?

OpenStudy (mathmale):

-60a - 6b = 6 => -10a - b = 1 -32a - 4b = 0 => -8a - b =0 To answer your question (to be continued)

OpenStudy (mathmale):

You could, if you wanted, divide the first equation by 60 or -60 or whatever, but subtracting 60 or 32 on both sides would be a mistake.

OpenStudy (anonymous):

Oh! I didn't even realize that! Wow, that was literally right in front of me. Okay! So now we have them reduced.

OpenStudy (mathmale):

Simply change all the signs in one of the equations on the right: -60a - 6b = 6 => -10a - b = 1 => 10a + b = -1 -32a - 4b = 0 => -8a - b =0 => -8a -b = 0

OpenStudy (mathmale):

Yes, you could reduce them if you want. But please take a look at this latest set of simultaneous equations. Do you see what to do next? If so, do it, please.

OpenStudy (mathmale):

This is "elimination by subtraction."

OpenStudy (anonymous):

SO 2a = 1 ?

OpenStudy (mathmale):

Check that. One small error.

OpenStudy (anonymous):

hmm. 10a + b = - 1

OpenStudy (anonymous):

so it would be 2a = -1

OpenStudy (mathmale):

2a = 1 should be 2a = -1. Agreed?

OpenStudy (mathmale):

yes, you had it correct before my question got posted!

OpenStudy (mathmale):

solve that for a, please

OpenStudy (anonymous):

a = -1/2

OpenStudy (mathmale):

Perfect. As I said earlier, I'd already solved this system using matrices on my TI-83 calculator. I also got a=-1/2. Nice work! Now go back to:

OpenStudy (mathmale):

10a + b = -1 and substitute a=-1/2, to find b. b=?

OpenStudy (anonymous):

b = 4

OpenStudy (mathmale):

Perfect! Now with b=4 and a=-1/2, go back to any one of the 3 original equations and find c.

OpenStudy (anonymous):

would c = -3 ?

OpenStudy (mathmale):

Again, right on! So, your quadratic becomes y=-0.5x^2 + 4x - 3

OpenStudy (mathmale):

going back to the original question, where you were given 3 points, you'll see that 2 of the points were (2,3) and (6,3). Look at the graph. Do these points seem to satisfy the equation (do they lie on the graph)?

OpenStudy (anonymous):

Awesome! And all of that only took up half of a page! I really wish I had a graphing calculator!

OpenStudy (anonymous):

And yes, they do fit the graph!

OpenStudy (mathmale):

yes, I'd recommend you get one if you're taking much more math. We could talk about this later. So you see, you have solved this problem completely. Starting from 3 points,

OpenStudy (mathmale):

you've come up with the equation of a quadratic function satisfied by the points.

OpenStudy (anonymous):

Okay! so now I have that, I can do it and I took good notes! I am going to show you some other things that have confused me if that is alright! I can finish the rest of these later

OpenStudy (mathmale):

Hope you have some way of capturing what we've typed here; it would be tremendously useful to you to have something to review. I know from experience that once I've learned just so much, my mind cannot absorb any more for a while, at least.

OpenStudy (mathmale):

Great minds think alike, as they say. :)

OpenStudy (mathmale):

Ready, aim, fire questions. Then choose a new problem on which to focus.

OpenStudy (mathmale):

got your photo; waiting for your questions.

OpenStudy (anonymous):

This is the question, there are two like this, I don't quite understand.

OpenStudy (anonymous):

I have to find a quadratic, then I have to estimate how much the cost would be for 24 in. x 36 in. photo

OpenStudy (mathmale):

image very slow in loading; let me check it again

OpenStudy (mathmale):

got image, looking at it now

OpenStudy (mathmale):

If I read this correctly, you'll need to average 24 and 36. Bet you've done that before.

OpenStudy (anonymous):

average them? Yes.

OpenStudy (mathmale):

my comment was out of order; sorry. Yes, you're right, we have to come up with a quadratic model just as we did in the last problem.

OpenStudy (anonymous):

If we average them, the mean would be 30

OpenStudy (mathmale):

Your table shows a bunch of photo sizes. Find the average of each set of dimensions. For exampel, for the 8 by 10, the average of the 2 given dimensions is (8+10)/2=9. Please go ahead and find the averages of the other sets of dimensions. We'll need this for ffidning the quadratic model.

OpenStudy (anonymous):

Oh! okay! yes I will do that. And I have a question. When we find the averages, do we have to use every plot point or just three like we did the last?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I'd suggest you do it for all of the table. we won't need that may points, however. As before, to find a quadratic model, we'll need just 3 points (you're right).

OpenStudy (mathmale):

Let's compare results. I've obtained: 9 10 16 16 18 19 27 27 36 39

OpenStudy (mathmale):

that was very insightful of you.

OpenStudy (anonymous):

Yes! that is exactly what i have, I have already started working on finding the quadratic. I am about to solve for a

OpenStudy (mathmale):

Yes, use the same quadratic model you started with before: y=ax^2+bx+c. choose any three of the points listed and write 3 equations, just as you did with the last problem.

OpenStudy (mathmale):

Once you have those 3 equations, you can solve them in the same way you did the previous system of linear equations. I'd prefer to move on to a new problem, but would be glad to go through the solution of this present system of linear equations if you like.

OpenStudy (mathmale):

Excuse me for a moment; I need to make a brief phone call. I'll be back!

OpenStudy (mathmale):

You don't have to type your 3 equations here unless you want to go through the solution with me.

OpenStudy (anonymous):

I am going to solve it and show you, I have a good feeling about this!

OpenStudy (mathmale):

OK. Just wanted to let you know that I'm back. Also delighted that you have a good feeling about your ability to solve this kind of problem!

OpenStudy (anonymous):

Okay, so I am trying to solve, but I have hit a problem. I subtracted the first and second equations and the first and third equations. But now I am stuck at: -156a - 6b = -7 -224a - 8b = -10

OpenStudy (mathmale):

I don't know whether or not that is correct as I haven't attempted the problem myself, but your results here do look appropriate.

OpenStudy (anonymous):

would you like me to take a picture of my work?

OpenStudy (mathmale):

It will be a little harder to solve this system because the 6 and the 8 coefficients of b are different; do you know what to do in a case like this or should we discuss what to do?

OpenStudy (anonymous):

I would like to discuss it please! Because I have no clue of what to do.

OpenStudy (mathmale):

Coming up.

OpenStudy (anonymous):

Would we isolate "a" ?

OpenStudy (mathmale):

In -156a - 6b = -7 -224a - 8b = -10 we want to eliminate either a or b, right? b would be easier to elim.

OpenStudy (mathmale):

Yes, elim. b would give us an expression in a which would be easy to solve for a. Let's keep the 1st eq'n as is. We want to change the 2nd eq'n so that the coeff. of b is +6; then -6b + 6b = 0 as desired. Right?

OpenStudy (anonymous):

Yes!

OpenStudy (mathmale):

Here's what I propose. Just multiply everythign in the 2nd row by (-6/8). Just do it, and then perhaps the reason for doing this will become apparent.

OpenStudy (anonymous):

Okay, would the first equation now become -156a - b = -1 ?

OpenStudy (anonymous):

And what do you mean by "multiply everything in the 2nd row by (-6/8)"?

OpenStudy (mathmale):

Okay, would the first equation now become -156a - b = -1 ? NO. The second row is -224a - 8b = -10. Multiplying everything in this row by (-6/8) results in 168a + 6b = 7.5. Please look this over and see whether you can agree with that or not.

OpenStudy (anonymous):

So where did the (-6/8) come from exactly?

OpenStudy (mathmale):

Now, the first row is still -156a - 6b = -7 I asked you to leave that as it is. The second row (new) is 168a + 6b = 7.5 Is that what y ou got? If so, add the like terms together (which will eliminate b). Mind if we finish this first, and then I'll answer your qu. about where that -6/8 came from.

OpenStudy (anonymous):

Yes! okay, so now I have 12a = 1/2